Circular No. 8439

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 TU_{12} (SIDING SPRING) G. Masi, Ceccano, Italy, reports further on their CCD images of this

comet (cf. IAUC 8436), noting that the tail was brighter close to the comet’s head on Nov. 12.0 UT; on Nov. 14.2 and 15.1 the tail was faint (barely visible) close to the nuclear condensation while peaking in intensity ∼ 100′′ and ∼ 130′′, respectively, from the comet’s head. On Nov. 14.2 and 15.1, the total tail length was > 5′.8 and > 10′, respectively, always pointing toward p.a. 70^{o}. Images taken on Nov. 17.04–17.07 show the tail to be definitely fainter than on the previous night, being barely visible for the first 1′ from the comet’s head but clearly brighter beyond that point (visible to > 6′, beyond the edge of the frames); the first 3′.3 of the tail was in p.a. 72^{o}, with a sudden change at that point to p.a. 73^{o}.

SUPERNOVAE 2004ga AND 2004gb A. Garg, Department of Physics, Harvard University; with S. Blondin,

P. Challis, C. Stubbs, and the SuperMACHO collaboration (cf. IAUC 7791; plus C. Aguilera, J. L. Prieto, D. Bramich, M. Huber, and D. Minniti, minus Hiriart, Lepischak, and Schmidt), announce the detection with the Cerro Tololo 4-m telescope (+ MOSAIC imager) of two supernovae located behind the Large Magellanic Cloud (the ‘VR’ magnitudes given below being their peak brightnesses); 4-m-telescope data of SN 2004ga were also obtained on Sept. 18, Oct. 6, 9, 11, 17, and 21, and of SN 2004gb also on Oct. 5, 9, 11, 13, 19, 21, and Nov. 6. Spectroscopy obtained at the Magellan Clay telescope (+ LDSS-2 spectrograph) on Nov. 3 and 5 indicates that SN 2004gb is of type Ia, ≈ 10 days after maximum, while SN 2004ga in a face-on galaxy is of type II; Magellan ‘MagIC’ observations of SN 2004gb were also obtained on Nov. 3, 5, and 6.

SN  2004 UT α_{2000}  δ_{2000}  Mag. z Offset
2004ga Sept. 11 5^{h}58^{m}11.^{s}5 −71^{o}37′00′′.3 19.7 0.11 238′′ W, 3′′ N
2004gb Oct. 15 4 55 22.3 −67 30 44.3 20.5 0.22  —

SUPERNOVA 2004fv IN NGC 6492 K. Krisciunas, University of Notre Dame, reports that spectra of SN

2004fv (IAUC 8430), taken by A. Pasten with the Cerro Tololo 1.5-m telescope on Nov. 16.0 UT and reduced by M. Salvo, show it to be a normal type-Ia supernova, ∼ 2 weeks after maximum light.

2004 November 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green